Fermi National Laboratory
 
Operation Summary for April 9, 2012 - April 11, 2012
 

Monday April 9

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, and SeaQuest experiments taking beam.

During the day shift, at 8:33 AM, the NuMI beam permit tripped off for the first of two times in five and a half hours due to a power supply’s (E:V118) output drifting high; everything reset. Operations established beam to the Meson FTBF T-992 experiment at 10:33 AM. At 12:50 PM, the MI RF South Anode Power Supply tripped off for the first of two times in about an hour due to an interlock fault; it reset, but after the second trip all the MI RF stations tripped off; they all reset. A MI person began a slow extraction study at 1:14 PM. Operations established beam to MTA at 1:48 PM. A MI person began a g-2 (g minus 2) study at 1:49 PM. Muon Ring personnel began a study at 2:23 PM. Operators reported that a Booster person had been conducting a study throughout the shift.

During the evening shift, the Muon Ring study ended at 4:14 PM. The Booster study ended at 4:36 PM. Operations disabled the MTA beam at 4:55 PM. The MI studies ended at 6:01 PM. At 10:22 PM, the NuMI permit tripped off on a MI incomplete RF gap, but none of the MI RF stations had tripped off; everything reset. Operations disabled T-992 beam at 11:14 PM. Operators reported at 11:59 PM that they had completed a 750 KeV area tune and a 400 MeV area steer.

Tuesday April 10

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 12:44 AM, the NuMI beam permit tripped off due to excessive ripple on a power supply (HV101); everything reset. At 2:25 AM, the NuMI beam permit tripped off due to V118; everything reset.

During the day shift, a Muon Ring person began a study at 8:52 AM. Operations established beam to T-992 at 9:10 AM. At 9:22 AM, the NuMI beam permit tripped off for the first of four times in two hours and 20 minutes due to V118; everything reset. The MiniBooNE beam permit tripped off at 9:45 AM for the first of two times in two hours due to an enclosure air flow (E:12AAIR) being a bit high; everything reset. Booster personnel reported setting up a new phase lock module at 10:39 AM for a one-hour study. At 2:07 PM, the NuMI permit tripped off on a MI incomplete RF gap, but none of the MI RF stations had tripped off; everything reset. Operations disabled beam to T-992 at 2:12 PM. Operations established beam to MTA at 2:20 PM for about an hour of beam.

During the evening shift, at 4:40 PM, a Linac Klystron RF station (KRF6) tripped off for the first of six times in about three hours; everything reset. Operations disabled beam to MTA at 5:18 PM.

Wednesday April 11

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, and SeaQuest.

During the midnight shift, at 3:03 AM, the NuMI permit tripped off due V118; everything reset. A Muon Ring person began a beam study at 4 AM. Operations disabled beam to SeaQuest at 7:03 AM in preparation for an access into enclosure C. The Muon Ring study ended at 7:55 AM. Operators had the Neutrino Area N01 enclosure ready for access by 7:59 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with Booster, MI, and Muon Ring study plans for today.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/12/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory